# Document Valuation in LLM Summaries: A Cluster Shapley Approach

### v2

@@ Appendix G Computation time of the Clustering Step
To precisely measure the computational cost of individual iterations, we conducted a separate timing experiment using identical code implementation on an Intel i7-13900K processor. We performed 1,000,000 iterations and completed them within $1.53\times 10^{-3}$ second, yielding an average time of $1.53\times 10^{-9}$ second per iteration. While we acknowledge that measurements at such small time scales can be subject to system-level variations due to factors such as CPU scheduling, memory access patterns, and hardware-specific optimizations, these results consistently demonstrate that the iteration overhead is orders of magnitude smaller than the LLM operations in our proposed Cluster Shapley Algorithm Step 2 in $\S$4.3, which require approximately 3.5 seconds per subset evaluation. This substantial difference in time scales confirms that the computational cost of our adaptive clustering approach remains negligible in the overall Shapley value calculation pipeline.
Appendix H Robustness Checks and Extensions
@@
H.1 Robustness Check using Claude for Evaluation
@@
H.2 Performance under MAPE and MSE Metrics
For completeness, we evaluate the performance of the Cluster Shapley algorithm using alternative error metrics—MAPE and MSE—and compare it to other benchmark methods in Figure A13. For ease of comparison, we replicate the results from Figure 5 as panel (c) in Figure A13. As shown, Cluster Shapley consistently outperforms the benchmarks across all three metrics.
@@ H.2 Performance under MAPE and MSE Metrics
H.3 Cluster Shapley using the Standard DBSCAN
@@
H.4 Sample Splitting Robustness Check
@@
H.5 Cluster Shapley with Cluster-Level Approximation Algorithms
@@
For the standard Cluster Shapley algorithm, we use the implementation described in the main paper. For the Monte Carlo–based variant, we vary two hyperparameters: the clustering diameter $\epsilon\in[0.01,1.0]$ and the number of Monte Carlo permutations used in Step 2. We evaluate all combinations of these hyperparameters and report the efficient frontier—i.e., the lowest approximation error achieved for each level of computational cost (measured by the number of unique subset evaluations).

### v3

@@ C.5 Computation time of the Clustering Step
To precisely measure the computational cost of individual iterations, we conducted a separate timing experiment using identical code implementation on an Intel i7-13900K processor. We performed 1,000,000 iterations and completed them within $1.53\times 10^{-3}$ seconds, yielding an average time of $1.53\times 10^{-9}$ second per iteration. While we acknowledge that measurements at such small time scales can be subject to system-level variations due to factors such as CPU scheduling, memory access patterns, and hardware-specific optimizations, these results consistently demonstrate that the iteration overhead is orders of magnitude smaller than the LLM operations in our proposed Cluster Shapley Algorithm Step 2 in $\S$4.3, which require approximately 3.5 seconds per subset evaluation. This substantial difference in time scales confirms that the computational cost of our adaptive clustering approach remains negligible in the overall Shapley value calculation pipeline.
Appendix D Robustness Checks and Extensions
@@
D.1 Robustness Check using Claude for Evaluation
@@
D.2 Performance under MAPE and MSE Metrics
For completeness, we evaluate the performance of the Cluster Shapley algorithm using alternative error metrics-MAPE and MSE-and compare it to other benchmark methods in Figure A13. For ease of comparison, we replicate the results from Figure 5 as panel (c) in Figure A13. As shown, Cluster Shapley consistently outperforms the benchmarks across all three metrics.
@@ D.2 Performance under MAPE and MSE Metrics
D.3 Cluster Shapley using the Standard DBSCAN
@@
D.4 Sample Splitting Robustness Check
@@
D.5 Cluster Shapley with Cluster-Level Approximation Algorithms
@@
For the standard Cluster Shapley algorithm, we use the implementation described in the main paper. For the Monte Carlo–based variant, we vary two hyperparameters: the clustering diameter $\epsilon\in[0.01,1.0]$ and the number of Monte Carlo permutations used in Step 2. We evaluate all combinations of these hyperparameters and report the efficient frontier-i.e., the lowest approximation error achieved for each level of computational cost (measured by the number of unique subset evaluations).